OpenStudy (anonymous):

You invest $700 in an account that pays an interest rate of 6.5%, compounded continuously. Calculate the balance after 20 years. Round to the nearest hundredth.

jimthompson5910 (jim_thompson5910):

A= Pe^(r*t) A = 700*e^(0.065*20) A = ???

jimthompson5910 (jim_thompson5910):

you type in exactly what you see on line 2 to evaluate

jimthompson5910 (jim_thompson5910):

your calculator should have an 'e' button e is like pi, but instead e = 2.718....